# Can We Generate Images with CoT? Let’s Verify and Reinforce Image Generation Step by Step

### v1

@@ Abstract
Chain-of-Thought (CoT) reasoning has been extensively explored in large models to tackle complex understanding tasks. However, it still remains an open question whether such strategies can be applied to verifying and reinforcing image generation scenarios. In this paper, we provide the first comprehensive investigation of the potential of CoT reasoning to enhance autoregressive image generation. We focus on three techniques: scaling test-time computation for verification, aligning model preferences with Direct Preference Optimization (DPO), and integrating these techniques for complementary effects. Our results demonstrate that these approaches can be effectively adapted and combined to significantly improve image generation performance. Furthermore, given the pivotal role of reward models in our findings, we propose the Potential Assessment Reward Model (PARM) and PARM++, specialized for autoregressive image generation. PARM adaptively assesses each generation step through a potential assessment approach, merging the strengths of existing reward models, and PARM++ further introduces a reflection mechanism to self-correct the generated unsatisfactory image. Using our investigated reasoning strategies, we enhance a baseline model, Show-o, to achieve superior results, with a significant +24% improvement on the GenEval benchmark, surpassing Stable Diffusion 3 by +15%. We hope our study provides unique insights and paves a new path for integrating CoT reasoning with autoregressive image generation. Code and models are released at https://github.com/ZiyuGuo99/Image-Generation-CoT.
Figure 1: Can We Verify and Reinforce Image Generation with Chain-of-Thought (CoT) Reasoning Strategies? Given the success of mathematical CoT reasoning in LLMs [1, 2] and LMMs [3, 4] (Left), we provide the first investigation to comprehensively explore the potential of applying current reasoning techniques to autoregressive image generation (Right), including test-time verification and preference alignment, with two newly proposed specialized reward models, termed PARM and PARM++.
[…]
@@ 1 Introduction
To alleviate these issues, we propose a specialized reward model for autoregressive image generation, termed Potential Assessment Reward Model (PARM). PARM adaptively verifies the generation process step by step with three delicately designed tasks: 1) Judge which step is clear and convincing enough to be evaluated, given that most early-stage images are typically blurry; 2) Assess whether the current step has the potential to yield a high-quality final image, since later-stage images generally do not change too much; and 3) Score the remaining final paths for selecting the best one, similar to an ORM. In this way, PARM can adaptively conduct assessment at appropriate steps (overcoming PRM’s scoring challenges), while effectively capturing fine-grained step-by-step cues (complementing the coarse ORM). Our experiments showcase that PARM significantly outperforms both ORM and PRM, which finally improves the baseline model by +24% on GenEval, as visualized in Figure 2, surpassing the advanced Stable Diffusion 3 [41] by +15%.
@@
  • We investigate specific adaption methods of techniques, including test-time verification, preference alignment, and ranking data curation, to autoregressive image generation, indicating their performance and complementarity.
@@ 2.2 ORM vs PRM as Test-time Verifiers
  • Zero-shot ORM: We leverage a pre-trained LLaVA-OneVision (7B) [39], an LMM with superior generalization, as our zero-shot ORM. We input the text prompt along with the generated image into LLaVA-OneVision, and devise a prompt template (detailed in the Appendix) to activate its visual understanding capabilities. This model assesses the quality of candidate images, providing binary responses, ‘yes’ (good quality) or ‘no’ (low quality). The candidate image with the highest probability of ‘yes’ is then selected as the final output.
  • ORM Ranking Data Curation: To enhance the accuracy of outcome rewards, we curate a dataset of 288K text-to-image ranking examples for fine-tuning ORM. First, we prompt GPT-4 [47] to generate a list of 200 countable daily object names with specific colors. Using these objects, we apply the six object-centered prompt templates from GenEval, constructing a diverse set of 13K text prompts. We perform a strict filtering to ensure that these prompts do not overlap with the GenEval test samples. Then, using our baseline model, Show-o, we synthesize around 50 images per prompt at a high temperature. After that, we label each image with a binary annotation of ‘yes’ or ‘no’ using the evaluation metric in GenEval.
PRM.
Different from ORM that evaluates only the final output, we utilize PRM to provide a reward score to each candidate with different steps throughout the generation process. Similar to our previous investigation, we start from a zero-shot PRM, LLaVA-OneVision, and then curate 10K step-wise text-to-image ranking data to obtain a fine-tuned PRM. We refer to the Appendix for detailed implementation of PRM.
Experiments and Insights.
@@ 2.4 DPO Alignment plus Test-time Verifiers
Table 3: Performance Comparison on the GenEval [38] Benchmark. Compared to existing diffusion and autoregressive models, we investigate the potential of Chain-of-Thought (CoT) reasoning strategies in text-to-image generation. ‘Ft. ORM’ and ‘It. DPO’ denote the fine-tuned ORM and iterative DPO [46]. PARM refers to our Potential Assessment Reward Model specialized for autoregressive image generation. We adopt the best-of-$20$ selection for test-time verifiers, highlighting our best result in green and the previous best model in red.
Model Test-time Verifier Preference Alignment Reward Guidance Single object Two object Counting Colors Position Attribute binding Overall PixArt-$\alpha$ [53] - - - 0.98 0.50 0.44 0.80 0.08 0.07 0.48 SD v2.1 [54] - - - 0.98 0.51 0.44 0.85 0.07 0.17 0.50 DALL-E 2 [43] - - - 0.94 0.66 0.49 0.77 0.10 0.19 0.52 SDXL [44] - - - 0.98 0.74 0.39 0.85 0.15 0.23 0.55 SD 3 (d=24) [41] - - - 0.98 0.74 0.63 0.67 0.34 0.36 0.62 LlamaGen [30] - - - 0.71 0.34 0.21 0.58 0.07 0.04 0.32 Chameleon [55] - - - - - - - - - 0.39 LWM [56] - - - 0.93 0.41 0.46 0.79 0.09 0.15 0.47 SEED-X [57] - - - 0.97 0.58 0.26 0.80 0.19 0.14 0.49 Show-o [5] - - - 0.95 0.52 0.49 0.82 0.11 0.28 0.53 Ft. ORM - - 0.99 0.72 0.65 0.84 0.25 0.33 0.63 - It. DPO - 0.98 0.72 0.53 0.84 0.40 0.46 0.65 Ft. ORM It. DPO Ft. ORM 0.98 0.84 0.64 0.85 0.66 0.52 0.75 PARM - - 0.99 0.77 0.68 0.86 0.29 0.45 0.67 - It. DPO PARM 0.97 0.75 0.60 0.83 0.54 0.53 0.69 PARM It. DPO - 0.98 0.83 0.64 0.84 0.59 0.62 0.74 PARM It. DPO PARM 0.99 0.86 0.67 0.84 0.66 0.64 0.77
@@ 3 Potential Assessment Reward Model
To empower PARM with robust capabilities, we curate a new ranking dataset with 400K instances by re-annotating the 13K text prompts from ORM ranking data. The dataset is structured into three subsets corresponding to the three evaluation tasks, containing 120K, 80K, and 200K instances, respectively. Please refer to the Appendix for detailed data formats.
@@
  • With PARM, our baseline model (Show-o) is enhanced to achieve leading generation performance. Compared to other image generation models in Table 3, our best-performing configuration, i.e., integrating PARM with iterative DPO in both post-training and test-time, achieves a score of 77%, improving the baseline by +24% and surpassing the advanced Stable Diffusion 3 [41] by +15%. In particular, substantial gains are observed in ‘Two Obj.’, ‘Colors’, ‘Position’, and ‘Attribute binding’ emphasizing the robustness in handling challenging compositional generation. In Figures 7, 8, 9, 10, and 11, we showcase extensive qualitative examples. We observe that the baseline model often generates inaccurate spatial relationships, strange appearances, and object attributes. In contrast, our approach consistently mitigates such issues, ensuring that the spatial relations, object features, and overall fidelity to the text prompt are preserved.
@@ 5 Conclusion
Appendix A Related Work
Scaling Test-time Computation.
Reinforced Preference Alignment.
Autoregressive Image Generation.
Appendix B Data and Implementation Details
B.1 ORM
Zero-shot ORM.
[…]
ORM Ranking Data Curation.
[…]
B.2 PRM
Zero-shot PRM.
[…]
PRM Ranking Data Curation.
[…]
Fine-tuned PRM.
With the step-wise ranking data, the LLaVA-OneVision is fine-tuned to boost the visual comprehension of intermediate-step images. The data format and training configurations are the same as those used for fine-tuning the ORM. After training, the PRM becomes more capable of interpreting blurry images within the decoding process for more accurate step-by-step selection.
B.3 PARM
Clarity Judgment Data (120K).
Through comprehensive analysis, we observe that the baseline model (Show-o) typically produces its first clear image between steps 8 and 12 within the 18-step generation, qualifying it for potential assessment. Based on this, we simplify the annotation by labeling steps after 11 as ‘yes’ and those before 10 as ‘no’. Although this approach is static, the trained PARM acquires generalization skills to adaptively identify the first ‘yes’ label within steps 8$\sim$12 during inference. The data format is shown below:
[…]
Potential Assessment Data (80K).
[…]
Best-of-$N^{\prime}$ Selection Data (200K).
[…]
Appendix C Additional Results
Quantitative Results.
In Table 5, we present a comprehensive performance comparison on GenEval [38] between previous diffusion and autoregressive models, and Shwo-o equipped with our investigated reasoning strategies. Substantial improvement for text-to-image generation are observed using different reasoning techniques. With PARM, the gains in complex attributes, such as ‘Two Obj.’, ‘Counting’, ‘Position’, and ‘Attribute binding’ emphasize the robustness of our approach in handling challenging aspects of compositional generation, setting a new standard in text-to-image performance. In Figures 13 and 14, we present the performance of test-time verification integrated with DPO [46] and iterative DPO, respectively, instead of the test-time verification only in Figure 2 of the main paper. As shown, our propose PARM both achieves the best results as the $N$ increases for best-of-$N$ selection.

### v2

@@ Abstract
Chain-of-Thought (CoT) reasoning has been extensively explored in large models to tackle complex understanding tasks. However, it still remains an open question whether such strategies can be applied to verifying and reinforcing image generation scenarios. In this paper, we provide the first comprehensive investigation of the potential of CoT reasoning to enhance autoregressive image generation. We focus on three techniques: scaling test-time computation for verification, aligning model preferences with Direct Preference Optimization (DPO), and integrating these techniques for complementary effects. Our results demonstrate that these approaches can be effectively adapted and combined to significantly improve image generation performance. Furthermore, given the pivotal role of reward models in our findings, we propose the Potential Assessment Reward Model (PARM) and PARM++, specialized for autoregressive image generation. PARM adaptively assesses each generation step through a potential assessment approach, merging the strengths of existing reward models, and PARM++ further introduces a reflection mechanism to self-correct the generated unsatisfactory image, which is the first to incorporate reflection in autoregressive image generation. Using our investigated reasoning strategies, we enhance a baseline model, Show-o, to achieve superior results, with a significant +24% improvement on the GenEval benchmark, surpassing Stable Diffusion 3 by +15%. We hope our study provides unique insights and paves a new path for integrating CoT reasoning with autoregressive image generation. Code and models are released at https://github.com/ZiyuGuo99/Image-Generation-CoT.
keywords:
Autoregressive Image Generation, Chain-of-Thought (CoT) Reasoning, Scaling Test-time Computation, Preference Alignment, Self Reflection
Figure 1: Can We Verify and Reinforce Image Generation with Chain-of-Thought (CoT) Reasoning Strategies? Given the success of mathematical CoT reasoning in LLMs [1, 2] and LMMs [3, 4] (Left), we provide the first investigation to comprehensively explore the potential of applying current reasoning techniques to autoregressive image generation (Right), including test-time verification, preference alignment and reflection, with two newly proposed specialized reward models, termed Potential Assessment Reward Model (PARM) and PARM++.
@@ 1 Introduction
To alleviate these issues, we propose a specialized reward model for autoregressive image generation, termed Potential Assessment Reward Model (PARM). PARM adaptively verifies the generation process step by step with three delicately designed tasks: 1) Judge which step is clear and convincing enough to be evaluated, given that most early-stage images are typically blurry; 2) Assess whether the current step has the potential to yield a high-quality final image, since later-stage images generally do not change too much; and 3) Score the remaining final paths for selecting the best one, similar to an ORM. In this way, PARM can adaptively conduct assessment at appropriate steps (overcoming PRM’s scoring challenges), while effectively capturing fine-grained step-by-step cues (complementing the coarse ORM). Our experiments showcase that PARM significantly outperforms both ORM and PRM, which finally improves the baseline model by +24% on GenEval, as visualized in Figure 2, surpassing the advanced Stable Diffusion 3 [41] by +15%. To further harness CoT-style reasoning, we introduce a reflection mechanism in PARM++, enabling the model to self-evaluate and refine its outputs in a multi-step, introspective manner.
@@
  • We investigate specific adaption methods of techniques, including test-time verification, preference alignment, ranking data curation and reflection, to autoregressive image generation, indicating their performance and complementarity.
@@
2 Related Work
Scaling Test-time Computation.
Reinforced Preference Alignment.
Autoregressive Image Generation.
@@ 3.2 ORM vs PRM as Test-time Verifiers
[…]
PRM.
Different from ORM that evaluates only the final output, we utilize PRM to provide a reward score to each candidate with different steps throughout the generation process. Similar to our previous investigation, we start from a zero-shot PRM, LLaVA-OneVision, and then curate 10K step-wise text-to-image ranking data to obtain a fine-tuned PRM.
[…]
  • Fine-tuned PRM: With the step-wise ranking data, the LLaVA-OneVision is fine-tuned to boost the visual comprehension of intermediate-step images. The data format and training configurations are the same as those used for fine-tuning the ORM. After training, the PRM becomes more capable of interpreting blurry images within the decoding process for more accurate step-by-step selection.
Experiments and Insights.
@@ 4 Potential Assessment Reward Model
  • Clarity Judgment Data (120K): Through comprehensive analysis, we observe that the baseline model (Show-o) typically produces its first clear image between steps 8 and 12 within the 18-step generation, qualifying it for potential assessment. Based on this, we simplify the annotation by labeling steps after 11 as ‘yes’ and those before 10 as ‘no’. Although this approach is static, the trained PARM acquires generalization skills to adaptively identify the first ‘yes’ label within steps 8$\sim$12 during inference. The data format is shown below:
[…]
@@
Table 4: Generalization Results of DPO, Finetuned ORM, and PARM on Sequential Autoregressive Generative Models, LlamaGen [72] and Janus-Pro [90]. ‘Zs.’ and ‘Ft.’ denote the zero-shot, fine-tuned verifiers. PARM refers to our proposed Potential Assessment Reward Model specialized for autoregressive image generation. We adopt the best-of-$20$ selection for test-time verifiers. Our methods consistently improve performance across both models, highlighting their model-agnostic and plug-and-play nature.
Model Method Single object Two object Counting Colors Position Attribute binding Overall LlamaGen [72] - 0.71 0.34 0.21 0.58 0.07 0.04 0.32 Zs. ORM 0.75 0.37 0.24 0.61 0.08 0.06 0.35 Ft. ORM 0.80 0.41 0.29 0.65 0.10 0.08 0.39 DPO 0.78 0.43 0.30 0.63 0.19 0.14 0.41 PARM 0.84 0.45 0.31 0.67 0.20 0.21 0.46 Janus-Pro [90] - 0.99 0.89 0.59 0.90 0.79 0.66 0.80 Zs. ORM 0.99 0.91 0.77 0.90 0.79 0.67 0.84 Ft. ORM 0.99 0.93 0.82 0.92 0.91 0.76 0.89 DPO 0.99 0.89 0.65 0.92 0.82 0.72 0.83 PARM 1.00 0.95 0.80 0.93 0.91 0.85 0.91
@@
  • With PARM, our baseline model (Show-o) is enhanced to achieve leading generation performance. Compared to other image generation models in Table 3, our best-performing configuration, i.e., integrating PARM with iterative DPO in both post-training and test-time, achieves a score of 77%, improving the baseline by +24% and surpassing the advanced Stable Diffusion 3 [41] by +15%. In particular, substantial gains are observed in ‘Two Obj.’, ‘Colors’, ‘Position’, and ‘Attribute binding’ emphasizing the robustness in handling challenging compositional generation. In Figures 6 and 7, we present the performance of test-time verification integrated with DPO [62] and iterative DPO, respectively, instead of the test-time verification only in Figure 2. As shown, our propose PARM both achieves the best results as the $N$ increases for best-of-$N$ selection. Additionally, in Figures 10, 11, 12, 13, and 14, we showcase extensive qualitative examples. We observe that the baseline model often generates inaccurate spatial relationships, strange appearances, and object attributes. In contrast, our approach consistently mitigates such issues, ensuring that the spatial relations, object features, and overall fidelity to the text prompt are preserved.
  • Generalization to Sequential Autoregressive Generative Models. While our method is validated on Show-o, which uses adaptive generation ordering with a discrete diffusion approach, we further extend it to sequential autoregressive models that generate image tokens via standard next-token prediction. Specifically, we apply DPO, Finetuned ORM, and PARM to LlamaGen-3B [72] and Janus-Pro-7B [90], both of which follow token-by-token generation in a fixed order. As shown in Table 4, our method consistently improves performance across both models without requiring retraining or architectural changes, demonstrating its generality and plug-and-play nature.